# Ambiguous Online Learning

### v3

@@ 7 Apple Tasting and Additional Lower Bounds
There is a close connection between ambiguous online learning and so-called apple tasting (see Helmbold et al. (2000)). In apple tasting, there is a hypothesis class $\mathcal{H}\subseteq\{X\to\{0,1\}\}$ and the learner’s output is also in $\{0,1\}$. As in classical online learning, whenever the learner’s prediction differs from the true label, this counts as a mistake. In contrast to classical online learning, when the learner’s prediction is 0, the learner receives no feedback (whereas if the learner’s prediction is 11, the true label is revealed). We will use the notation $A:(X\times\{0,1\})^{*}\times X\to\{0,1\}$ for apple tasting learners, where “no feedback” is represented by 11.
@@ 8 Conclusion
Qualitatively, instead of the two possible behaviors of classical theory ($O(1)$ mistakes and NN mistakes), we have three possible behaviors, with the new option of $O(\sqrt{N})$ mistakes. The upper bound of this new option is attained using the WAA, an algorithm reminiscent of weighted majority. The coefficient depends both on the classical Littlestone dimension $\mathrm{L}_{\mathfrak{P}}(\mathcal{H})$ and the combinatorial invariant $\underline{\mathrm{PD}}(\mathcal{H})$. As in classical theory, there is at most a constant (w.r.t. NN) gap between the mistake bounds of deterministic and randomized learners. However, our current best characterization of this gap involves yet another new combinatorial invariant, namely $\overline{\mathrm{PD}}(\mathcal{H})$. There are many natural questions for further study:
@@ Acknowledgments
This work was supported by the Machine Intelligence Research Institute in Berkeley, California, the Effective Ventures Foundation USA in San Francisco, California, the Advanced Research + Invention Agency (ARIA) of the United Kingdom, and the Survival and Flourishing Corporation. The author wishes to thank Shay Moran for insightful conversations, especially the suggestions to look into apple tasting and to mimic reduction to finite classes in nonrealizable online learning. She also thanks Alexander Appel and Marcus Ogren for reading drafts, pointing our errors and providing useful suggestions.
@@ Appendix A Properties of Combinatorial Invariants
To prove Proposition 3.4, we traverse a path going down from the root. At each vertex $v$ on the path, there will be a leaf label $h$ for which the $E_{0}$-edge of $v$ is relevant, and we’$l$ continue the path through one of the other edges which possesses an edge label outside of $h(\mathbf{x}_{v})$. The leaf labels we encounter in this way are all different.
of Proposition 3.4.
@@
The following simple lemmas will be useful to prove Proposition 3.8.
@@
Consider any $Y$-lattice $\Lambda$. For any $A,B\subseteq Y$, if $A\subseteq B$ then $A^{(\Lambda)}\subseteq B^{(\Lambda)}$.
@@
Notice that $A\subseteq B\subseteq B^{(\Lambda)}$. Also, $B^{(\Lambda)}\in\Lambda$. Hence, $A^{(\Lambda)}\subseteq B^{(\Lambda)}$. ∎
Lemma A.5.
Consider any $Y$-lattice $\Lambda$. For any $A,B\subseteq Y$:
$(A\cup B)^{(\Lambda)}=(A^{(\Lambda)}\cup B^{(\Lambda)})^{(\Lambda)}$
Proof.
Notice that $A\subseteq A\cup B$ and hence by Lemma A.4, $A^{(\Lambda)}\subseteq(A\cup B)^{(\Lambda)}$. Similarly, $B^{(\Lambda)}\subseteq(A\cup B)^{(\Lambda)}$. Therefore,
$A^{(\Lambda)}\cup B^{(\Lambda)}\subseteq(A\cup B)^{(\Lambda)}$
Since $(A\cup B)^{(\Lambda)}\in\Lambda$, this implies
$(A^{(\Lambda)}\cup B^{(\Lambda)})^{(\Lambda)}\subseteq(A\cup B)^{(\Lambda)}$
Conversely, $A\cup B\subseteq A^{(\Lambda)}\cup B^{(\Lambda)}$ and by Lemma A.4
$(A\cup B)^{(\Lambda)}\subseteq(A^{(\Lambda)}\cup B^{(\Lambda)})^{(\Lambda)}$
Proposition 3.8 is proved by showing that whenever $B$ is minimal s.t. $A\subseteq B^{(\Lambda)}$, every $C\subseteq B$ has a unique $\Lambda$-hull.
of Proposition 3.8.
Consider any $A\subseteq Y$. Let $B\subseteq A$ be of minimal cardinality such that $A\subseteq B^{(\Lambda)}$. Consider any $C_{1},C_{2}\subseteq B$ s.t. $C_{1}\neq C_{2}$. Assume w.l.o.g. that $C_{1}\setminus C_{2}\neq\emptyset$.
We claim that $C_{1}^{(\Lambda)}\setminus C_{2}^{(\Lambda)}\neq\emptyset$. Indeed, assume to the contrary that $C_{1}^{(\Lambda)}\subseteq C_{2}^{(\Lambda)}$. Denote $B_{1}:=B\setminus(C_{1}\setminus C_{2})$. Using Lemma A.5 twice, we get
$B_{1}^{(\Lambda)}$ $=(B_{1}\cup C_{2})^{(\Lambda)}$
$=(B_{1}^{(\Lambda)}\cup C_{2}^{(\Lambda)})^{(\Lambda)}$
$=(B_{1}^{(\Lambda)}\cup C_{1}^{(\Lambda)}\cup C_{2}^{(\Lambda)})^{(\Lambda)}$
$=(B_{1}\cup C_{1}\cup C_{2})^{(\Lambda)}$
$=B^{(\Lambda)}$
This contradicts the assumption that $B$ is of minimal cardinality.
Consider the mapping $f:2^{B}\to\Lambda$ given by $f(C):=C^{(\Lambda)}$. By the above, $f$ is injective. Hence, $2^{|B|}\leq|\Lambda|$ and therefore $\mathcal{C}_{\Lambda}(A)\leq\log_{2}|\Lambda|$. Since this holds for any $A\subseteq Y$, we get $\overline{\mathrm{PD}}(\Lambda)\leq\log_{2}|\Lambda|$. ∎
In order to calculate $\underline{\mathrm{PD}}(\Lambda)$ in Example 3.9, we observe that every $A\in\Lambda$ is a hyperrectangle and the $\Lambda$-hull of its two extreme corners equals AA. In order to calculate $\overline{\mathrm{PD}}(\Lambda)$, for any $A\subseteq[n]^{d}$ we take, for each dimension, a point in AA with minimal coordinate along this dimension and a point in AA with maximal coordinate. These are $2d$ points whose $\Lambda$-hull is AA.
of Example 3.9.
@@
Let $B_{ab}:=\{a,b\}$. Then, $B_{ab}^{((\Lambda_{d,n}))}=A_{ab}$. Hence, $\mathcal{C}_{\Lambda_{d,n}}(A_{ab})\leq 2$ and $\underline{\mathrm{PD}}(\Lambda_{d,n})\leq 2$.
For any $B\subseteq Y$ s.t. $|B|=1$, we have $B\in\Lambda_{d,n}$ and hence $B^{(\Lambda)}=B$. Therefore, for $a\neq b$, $\mathcal{C}_{\Lambda_{d,n}}(A_{ab})=2$ and hence $\underline{\mathrm{PD}}(\Lambda_{d,n})=2$.
@@
Then, $A\subseteq B^{((\Lambda_{d,n}))}$ and $|B|\leq 2d$. Hence, $\mathcal{C}_{\Lambda_{d,n}}(A)\leq 2d$ and $\overline{\mathrm{PD}}(\Lambda_{d,n})\leq 2d$.
@@
Then, for any $A\subseteq A^{*}$, $A^{((\Lambda_{d,n}))}\cap A^{*}=A$. Hence, $\mathcal{C}_{\Lambda_{d,n}}((A^{*})^{(|\Lambda_{d,n}))})\geq|A^{*}|=2d$ and so $\overline{\mathrm{PD}}(\Lambda_{d,n})=2d$. ∎
@@
of Proposition 3.10.
Let $T_{0}=(V,v_{0},P,\mathbf{x},\mathbf{y},\mathbf{h},E_{0})$ be an $\mathcal{H}$-tree s.t. $\operatorname{rank}(T_{0})\geq r$ and $\operatorname{dep}(T_{0})\leq n$. Suppose that $v_{1}\in V$ is of arity $m>\overline{\mathrm{PD}}(\mathcal{H})+1$. That is, $|P^{-1}(v_{1})|=m$. Define
@@
There exists $B\subseteq A$ s.t. $A\subseteq B^{((\Lambda(\mathcal{H})))}$ and $|B|\leq\overline{\mathrm{PD}}(\mathcal{H})$. Define $F\subseteq E$ by
@@
Consider some $u\in V_{f}^{\prime}$ and denote $\alpha:=\mathbf{h}_{u}(\mathbf{x}_{v_{1}})$. Consider also some $b\in P^{\prime*}(u)$ s.t. $b\notin R_{T_{1}}(u)$. If $P^{\prime}(b)\neq v_{1}$ then clearly $b\notin R_{T_{0}}(u)$. If $P^{\prime}(b)=v_{1}$ then for all $a\in F_{1}\setminus\{a_{1}\}$ we have $\mathbf{y}_{a}\in\alpha$, implying that $B\subseteq\alpha$. It follows that $B^{((\Lambda(\mathcal{H})))}\subseteq\alpha$ and hence $A\subseteq\alpha$. Therefore, $b\notin R_{T_{0}}(u)$ in this case as well. We conclude that $R_{T_{0}}(u)\subseteq R_{T_{1}}(u)$ (in fact they are equal).
Therefore, $\operatorname{rank}(T_{1})\geq\operatorname{rank}(T_{0})$. Obviously, also $\operatorname{dep}(T_{1})\leq\operatorname{dep}(T_{0})$. At the same time, the arity of $v_{1}$ in $T_{1}$ is $|F_{1}|\leq\overline{\mathrm{PD}}(\mathcal{H})+1$. This construction can be repeated until we get the desired $T$. ∎
@@
Then, there exists $a\in P^{-1}(v^{*})$ s.t. $\mathbf{y}_{a}\notin A^{((\Lambda(\mathcal{H})))}$.
@@
Let $a^{*}\in E_{0}$ be s.t. $P(a^{*})=v^{*}$. Since $a^{*}$ is non-redundant, there exists $u\in V_{f}$ and $a\in E$ s.t. $a^{*}\in P^{*}(u)$, $P(a)=v^{*}$ and $\mathbf{y}_{a}\notin\mathbf{h}_{u}(x^{*})$. Denote $y^{*}:=\mathbf{y}_{a}$ and $B:=\mathbf{h}_{u}(x^{*})$. For any $b\in P^{*}(v^{*})\setminus\{v_{0}\}$ s.t. $\mathbf{x}_{P(b)}=x^{*}$, we have $b\in P^{*}(u)$ and therefore $\mathbf{y}_{b}\in B$. We conclude that $A\subseteq B$. Since $B\in\Lambda(\mathcal{H})$, it follows that $A^{((\Lambda(\mathcal{H})))}\subseteq B$. On the other hand $y^{*}\notin B$, implying that $y^{*}\notin A^{((\Lambda(\mathcal{H})))}$. ∎
@@
Choose $v_{k+1}\in V$ to be s.t. $P(v_{k+1})=v_{k}$ and $\mathbf{y}_{v_{k+1}}\notin A_{k}^{((\Lambda(\mathcal{H})))}$. By Lemma A.6, this is possible as long as $v_{k}\notin V_{f}$ (hence $v_{n}\in V_{f}$).
@@
For any $k<n$, denote $B_{k}:=A_{k}^{((\Lambda(\mathcal{H})))}$. Notice that whenever $i,j\in I_{x}$ and $i<j$, we have $\mathbf{y}_{v_{i+1}}\notin B_{i}$, $\mathbf{y}_{v_{i+1}}\in A_{j}$ and therefore $B_{j}\setminus B_{i}\neq\emptyset$. Hence, $B_{i}\subset B_{j}$.
Let $k_{x}:=\max I_{x}$ and define $A_{(x)}:=A_{k_{x}}\cup\{\mathbf{y}_{v_{k_{x}+1}}\}$ and $B_{(x)}:=A_{(x)}^{((\Lambda(\mathcal{H})))}$. Then, $\mathbf{y}_{v_{k_{x}+1}}\notin B_{k_{x}}$, $\mathbf{y}_{v_{k_{x}+1}}\in A_{(x)}$ and hence $B_{k_{x}}\subset B_{(x)}$. We constructed a chain of length $|I_{x}|+1$ and hence $|I_{x}|\leq l-1$. Since $n]=\bigcup_{x}I_{x}$, we get $n\leq l|X|$. Therefore $|R_{T}(v_{n})|\leq n\leq l|X|$ and $\operatorname{rank}(T)\leq l|X|$. ∎

### v4

@@ 7 Apple Tasting and Additional Lower Bounds
There is a close connection between ambiguous online learning and so-called apple tasting (see Helmbold et al. (2000)). In apple tasting, there is a hypothesis class $\mathcal{H}\subseteq\{X\to\{0,1\}\}$ and the learner’s output is also in $\{0,1\}$. As in classical online learning, whenever the learner’s prediction differs from the true label, this counts as a mistake. In contrast to classical online learning, when the learner’s prediction is 0, the learner receives no feedback (whereas if the learner’s prediction is $1$, the true label is revealed). We will use the notation $A:(X\times\{0,1\})^{*}\times X\to\{0,1\}$ for apple tasting learners, where “no feedback” is represented by $1$.
@@ 8 Conclusion
Qualitatively, instead of the two possible behaviors of classical theory ($O(1)$ mistakes and $N$ mistakes), we have three possible behaviors, with the new option of $O(\sqrt{N})$ mistakes. The upper bound of this new option is attained using the WAA, an algorithm reminiscent of weighted majority. The coefficient depends both on the classical Littlestone dimension $\mathrm{L}_{\mathfrak{P}}(\mathcal{H})$ and the combinatorial invariant $\mathrm{PD}(\mathcal{H})$. As in classical theory, there is at most a constant (w.r.t. $N$) gap between the mistake bounds of deterministic and randomized learners. Our current best characterization of this gap is in terms of $\mathrm{VC}(\Lambda(\mathcal{H}))$. There are many natural questions for further study:
@@ Acknowledgments
This work was supported by the Machine Intelligence Research Institute in Berkeley, California, the Effective Ventures Foundation USA in San Francisco, California, the Advanced Research + Invention Agency (ARIA) of the United Kingdom, and the Survival and Flourishing Corporation. The author wishes to thank Shay Moran for insightful conversations, especially the suggestions to look into apple tasting and to mimic reduction to finite classes in nonrealizable online learning. She also wishes to thank Bogdan Chornomaz for pointing out the relation between VC dimension and $\Lambda$-complexity (Lemma A.4). Additionally, she thanks Alexander Appel and Marcus Ogren for reading drafts, pointing out errors, and providing useful suggestions.
@@ Appendix A Properties of Combinatorial Invariants
To prove Proposition 3.4, we traverse a path going down from the root. At each vertex $v$ on the path, there will be a leaf label $h$ for which the $E_{0}$-edge of $v$ is relevant, and we’ll continue the path through one of the other edges which possesses an edge label outside of $h(\mathbf{x}_{v})$. The leaf labels we encounter in this way are all different.
Proof of Proposition 3.4.
@@
The significance of $\mathrm{VC}(\Lambda)$ in our context comes from the following lemma.
@@
For any $Y$-lattice $\Lambda$, we have
$\mathrm{VC}(\Lambda)=\max_{A\subseteq Y}\mathcal{C}_{\Lambda}(A)$
@@
Consider any $\Lambda$-shattered set $A\subseteq Y$. Then, for any $B\subseteq A$, there is some $C\in\Lambda$ s.t. $B=A\cap C$. Therefore, $B^{(\Lambda)}\cap A=B$. It follows that $\mathcal{C}_{\Lambda}(A)=|A|$. Therefore,
$\mathrm{VC}(\Lambda)\leq\max_{A\subseteq Y}\mathcal{C}_{\Lambda}(A)$
Conversely, consider any $A\subseteq Y$ and let $k:=\mathcal{C}_{\Lambda}(A)$. Choose $B\subseteq A$ s.t. $|B|=k$ and $A\subseteq B^{(\Lambda)}$. Consider any $y\in B$ and denote $B_{y}:=B\setminus\{y\}$. Then, $y\not\in B_{y}^{(\Lambda)}$: otherwise we would have $B\subseteq B_{y}^{(\Lambda)}$ and therefore $A\subseteq B_{y}^{(\Lambda)}$, implying $\mathcal{C}_{\Lambda}(A)\leq k-1$. Consider any $C\subseteq B$ and define
$\hat{C}:=\bigcap_{y\in B\setminus C}B_{y}^{(\Lambda)}$
We get that $\hat{C}\in\Lambda$ and $C=B\cap\hat{C}$. Since this works for any $C$, we conclude that $B$ is a $\Lambda$-shattered set and hence $\mathrm{VC}(\Lambda)\geq k$. Therefore,
$\mathrm{VC}(\Lambda)\geq\max_{A\subseteq Y}\mathcal{C}_{\Lambda}(A)$
It is now trivial to prove Proposition 3.8.
Proof of Proposition 3.8.
By Lemma A.4,
$\mathrm{VC}(\Lambda)$ $=\max_{A\subseteq Y}\mathcal{C}_{\Lambda}(A)$
$\geq\max_{A\in\Lambda}\mathcal{C}_{\Lambda}(A)$
$=\mathrm{PD}(\Lambda)$
In order to calculate $\mathrm{PD}(\Lambda)$ in Example 3.9, we observe that every $A\in\Lambda$ is a hyperrectangle and the $\Lambda$-hull of its two extreme corners equals $A$. In order to calculate $\mathrm{VC}(\Lambda)$, for any $A\subseteq[n]^{d}$ we take, for each dimension, a point in $A$ with minimal coordinate along this dimension and a point in $A$ with maximal coordinate. These are $2d$ points whose $\Lambda$-hull is $A$.
Proof of Example 3.9.
@@
Let $B_{ab}:=\{a,b\}$. Then, $B_{ab}^{(\Lambda_{d,n})}=A_{ab}$. Hence, $\mathcal{C}_{\Lambda_{d,n}}(A_{ab})\leq 2$ and $\mathrm{PD}(\Lambda_{d,n})\leq 2$.
For any $B\subseteq Y$ s.t. $|B|=1$, we have $B\in\Lambda_{d,n}$ and hence $B^{(\Lambda)}=B$. Therefore, for $a\neq b$, $\mathcal{C}_{\Lambda_{d,n}}(A_{ab})=2$ and hence $\mathrm{PD}(\Lambda_{d,n})=2$.
@@
Then, $A\subseteq B^{(\Lambda_{d,n})}$ and $|B|\leq 2d$. Hence, $\mathcal{C}_{\Lambda_{d,n}}(A)\leq 2d$ and by Lemma A.4, $\mathrm{VC}(\Lambda_{d,n})\leq 2d$.
@@
Then, for any $A\subseteq A^{*}$, $A^{(\Lambda_{d,n})}\cap A^{*}=A$. Hence, $\mathcal{C}_{\Lambda_{d,n}}((A^{*})^{(\Lambda_{d,n})})\geq|A^{*}|=2d$ and by Lemma A.4, $\mathrm{VC}(\Lambda_{d,n})=2d$. ∎
@@
Proof of Proposition 3.10.
Let $T_{0}=(V,v_{0},P,\mathbf{x},\mathbf{y},\mathbf{h},E_{0})$ be an $\mathcal{H}$-tree s.t. $\operatorname{rank}(T_{0})\geq r$ and $\operatorname{dep}(T_{0})\leq n$. Suppose that $v_{1}\in V$ is of arity $m>\mathrm{VC}(\Lambda(\mathcal{H}))+1$. That is, $|P^{-1}(v_{1})|=m$. Define
@@
By Lemma A.4, there exists $B\subseteq A$ s.t. $A\subseteq B^{(\Lambda(\mathcal{H}))}$ and $|B|\leq\mathrm{VC}(\Lambda(\mathcal{H}))$. Define $F\subseteq E$ by
@@
Consider some $u\in V_{f}^{\prime}$ and denote $\alpha:=\mathbf{h}_{u}(\mathbf{x}_{v_{1}})$. Consider also some $b\in P^{\prime*}(u)$ s.t. $b\notin R_{T_{1}}(u)$. If $P^{\prime}(b)\neq v_{1}$ then clearly $b\notin R_{T_{0}}(u)$. If $P^{\prime}(b)=v_{1}$ then for all $a\in F_{1}\setminus\{a_{1}\}$ we have $\mathbf{y}_{a}\in\alpha$, implying that $B\subseteq\alpha$. It follows that $B^{(\Lambda(\mathcal{H}))}\subseteq\alpha$ and hence $A\subseteq\alpha$. Therefore, $b\notin R_{T_{0}}(u)$ in this case as well. We conclude that $R_{T_{0}}(u)\subseteq R_{T_{1}}(u)$ (in fact they are equal).
Therefore, $\operatorname{rank}(T_{1})\geq\operatorname{rank}(T_{0})$. Obviously, also $\operatorname{dep}(T_{1})\leq\operatorname{dep}(T_{0})$. At the same time, the arity of $v_{1}$ in $T_{1}$ is $|F_{1}|\leq\mathrm{VC}(\Lambda(\mathcal{H}))+1$. This construction can be repeated until we get the desired $T$. ∎
@@
Then, there exists $a\in P^{-1}(v^{*})$ s.t. $\mathbf{y}_{a}\notin A^{(\Lambda(\mathcal{H}))}$.
@@
Let $a^{*}\in E_{0}$ be s.t. $P(a^{*})=v^{*}$. Since $a^{*}$ is non-redundant, there exists $u\in V_{f}$ and $a\in E$ s.t. $a^{*}\in P^{*}(u)$, $P(a)=v^{*}$ and $\mathbf{y}_{a}\notin\mathbf{h}_{u}(x^{*})$. Denote $y^{*}:=\mathbf{y}_{a}$ and $B:=\mathbf{h}_{u}(x^{*})$. For any $b\in P^{*}(v^{*})\setminus\{v_{0}\}$ s.t. $\mathbf{x}_{P(b)}=x^{*}$, we have $b\in P^{*}(u)$ and therefore $\mathbf{y}_{b}\in B$. We conclude that $A\subseteq B$. Since $B\in\Lambda(\mathcal{H})$, it follows that $A^{(\Lambda(\mathcal{H}))}\subseteq B$. On the other hand $y^{*}\notin B$, implying that $y^{*}\notin A^{(\Lambda(\mathcal{H}))}$. ∎
@@
Choose $v_{k+1}\in V$ to be s.t. $P(v_{k+1})=v_{k}$ and $\mathbf{y}_{v_{k+1}}\notin A_{k}^{(\Lambda(\mathcal{H}))}$. By Lemma A.5, this is possible as long as $v_{k}\notin V_{f}$ (hence $v_{n}\in V_{f}$).
@@
For any $k<n$, denote $B_{k}:=A_{k}^{(\Lambda(\mathcal{H}))}$. Notice that whenever $i,j\in I_{x}$ and $i<j$, we have $\mathbf{y}_{v_{i+1}}\notin B_{i}$, $\mathbf{y}_{v_{i+1}}\in A_{j}$ and therefore $B_{j}\setminus B_{i}\neq\emptyset$. Hence, $B_{i}\subset B_{j}$.
Let $k_{x}:=\max I_{x}$ and define $A_{(x)}:=A_{k_{x}}\cup\{\mathbf{y}_{v_{k_{x}+1}}\}$ and $B_{(x)}:=A_{(x)}^{(\Lambda(\mathcal{H}))}$. Then, $\mathbf{y}_{v_{k_{x}+1}}\notin B_{k_{x}}$, $\mathbf{y}_{v_{k_{x}+1}}\in A_{(x)}$ and hence $B_{k_{x}}\subset B_{(x)}$. We constructed a chain of length $|I_{x}|+1$ and hence $|I_{x}|\leq l-1$. Since $n]=\bigcup_{x}I_{x}$, we get $n\leq l|X|$. Therefore $|R_{T}(v_{n})|\leq n\leq l|X|$ and $\operatorname{rank}(T)\leq l|X|$. ∎